Evaluate integral of c (xy^4) ds where c is right half of the circle x^2 + y^2 =16 traced out in a counter clockwise direction parameterized using x=4cost, y=4sint


1
Expert's answer
2022-03-04T09:32:12-0500

We can parametrize "C" by


"\vec r=\langle4\cos t, 4\sin t\rangle, -\pi/2\le t\le \pi/2"

Then we have


"\vec r'=\langle-4\sin t, 4\cos t\rangle"

"|\vec r'|=\sqrt{(-4\sin t)^2+(4\cos t)^2}=4"

"\int_Cxy^4ds=\displaystyle\int_{-\pi/2}^{\pi/2}(4\cos t)(4\sin t)^4(4)dt"

"=4^6[\dfrac{\sin ^5t}{5}]\begin{matrix}
 \pi/2 \\
 -\pi/2
\end{matrix}=\dfrac{2^{13}}{5}"